OpenStudy (anonymous):

Explain how to factor the following trinomials forms: x² + bx + c and ax² + bx + c. Is there more than one way to factor this? Show your answer using both words and mathematical notation.

OpenStudy (anonymous):

x^2+6x+8=(x+2)(x+4) Factoring 〖ax〗^2+bx+c/ mn=ac, m+n=b so 3*2=ac,3+2=b/ 〖6x〗^2+5x+1